Shift report 20000823 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Aug 23 2000 - 08:10:14 EDT

  • Next message: Jens Jorgen Gaardhoje: "Agenda for BRAHMS collaboration meeting"

    ------------------------------------------------------------
    Date: Wed Aug 23 08:10:13 2000, Shift: 00:00-08:00
    Supervisor: ykl, with: yb
    Log book#4 pages 144 - 146
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam was lost due to link quench problem at 1:32.
    The newly stored beam lasted from 5:04 till 7:17. An
    increased background is suspected because the rate of 
    trigger 1 is almost equivalent to the rate of trigger 4.
    2. The run 2403, continued from the previous shift gave
    516K of Trigger 1 with D5 at 350A with polarity A. Run 2406
    gave 301K with D5 at 350A with polarity B.
    3. T1 and T3 are not turned on. No new detector problems.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Aug 23 2000 - 08:24:33 EDT